For Drosophila Melanogaster how to determine whether the loci are assorting independently (lets say we are talking about two traits, ebony body and curved wing)

Would you have to perform a Chi-square testing for it?
Answer #1

Answer:

First we need to know the observed numbers.

Then we have to do testcross. The expected Dihybrid testcross ratio 1:1:1:1.

Based on the values, the table is prepared as below:

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

WT wing and body

44

45

3

9

0.15

Scalloped wing and WT body

42

45

-3

9

0.15

WT wing and ebony body

48

45

-2

4

0.2

Scalloped wing and ebony body

46

45

2

4

0.2

Total

180

180

0.7

Chi-square value = 0.70

Degrees of freedom = Number of catergories – 1

Df = 4 -1= 3

Critical value at probability of 0.05 = 7.81 (standard value)

The chi-square value of 0.70 is less than critical value of 7.81. We can accept the null hypothesis and the two genes are assorted independently.

----------------------------------------

Answer:

First we need to know the observed numbers.

Then we have to do testcross. The expected Dihybrid testcross ratio 1:1:1:1.

Based on the values, the table is prepared as below:

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

WT wing and bristles

44

45

3

9

0.15

apterous wing and WT body

42

45

-3

9

0.15

WT wing and spinless bristles

48

45

-2

4

0.2

apterous wing and pinless bristles

46

45

2

4

0.2

Total

180

180

0.7

Chi-square value = 0.70

Degrees of freedom = Number of catergories – 1

Df = 4 -1= 3

Critical value at probability of 0.05 = 7.81 (standard value)

The chi-square value of 0.70 is less than critical value of 7.81. We can accept the null hypothesis and the two genes are assorted independently.